# Scaling Laws of Synthetic Data for Language Models

### v2

@@ 4 SynthLLM: Web-Scale Synthetic Data Generation
In this section, we present details of SynthLLM for synthetic data generation. For specific target domain (e.g., mathematical reasoning), we begin by filtering high-quality and domain-specific reference documents from Fineweb-Edu [40], an open repository of web data (Section 4.1). Next, using these filtered reference documents, we construct large-scale, diverse questions (or prompts) by prompting open-source LLMs through three distinct methods, each designed to progressively increase question diversity (Section 4.2). Finally, we produce corresponding answers (or responses) for the generated questions, again employing open-source LLMs (Section 4.3).
@@ 4.1 Reference Documents Filtering
Figure 3: Overview of reference documents filtering. The whole process contains two classifiers: 1) cold-start classifier trained on generated positive samples (documents) and randomly sampled negative examples; 2) fine-grained classifier trained on previously filtered documents and additional randomly sampled documents with extra ratings. We train fine-grained classifier for 2 iterations.
Cold-start Domain Classifier.
For a given target domain, we train a binary classifier to filter web documents from Fineweb-Edu. While negative examples (documents) can be randomly sampled, obtaining sufficient positive examples (documents) for each domain is more challenging. To address this, we use synthetic documents as positive examples. Specifically, we generate synthetic domain-specific documents based on syllabi from GLAN [24], which cover 126 distinct disciplines. Each syllabus is produced by GPT-4 "(almost) from scratch", with further details available in [24]. As illustrated in Figure 3, we first use these syllabi to prompt a large language model (LLM) to generate synthetic documents representative of the target domain (e.g., blog posts, exam papers, and textbooks). Once we obtain the synthetic positive examples and randomly sampled negative examples from Fineweb-Edu, we train a cold-start domain classifier. Applying this classifier yields an initial set of domain-relevant documents, $\mathcal{D}^{0}$, from Fineweb-Edu.
Fine-grained Domain Classifier.
The initial reference document set, $\mathcal{D}^{0}$, is typically small. To expand it, we iteratively classify Fineweb-Edu documents based on previously identified references, $\mathcal{D}^{t-1}$. At each iteration $t$, we randomly sample an additional subset $\mathcal{D}^{t-1}_{R}$ from Fineweb-Edu and instruct GPT-4o to evaluate all documents in $\mathcal{D}^{t-1}\cup\mathcal{D}^{t-1}_{R}$ on a 1–10 scale, considering relevance to the target domain, clarity, and language quality. These ratings are then used to train a fine-grained classifier, which we apply to Fineweb-Edu to construct the next iteration’s reference set, $\mathcal{D}^{t}$. We retain documents with scores above a threshold of 6.5, as it yields satisfactory performance in initial experiments. Empirically, two iterations suffice to generate a high-quality and sufficiently large reference corpus.
@@ 4.2 Document-Grounded Question Generation
After applying the filtering process described in Section 4.1, we obtain the filtered document set $\mathcal{D}$. We then generate questions through three distinct methods, each relying on either single-document or multi-document contexts, with the goal of progressively increasing question diversity. Let $\mathbf{M}^{Q}$ denote the LLM employed for question generation.
Level-1 Generator.
In the first method, we generate questions from a single reference document $d\in\mathcal{D}$ (where $\mathcal{D}$ is filtered document set). Because $d$ may already contain questions relevant to target domain, we first prompt the LLM $\mathbf{M}^{Q}$ to determine whether $d$ includes any existing questions. If such questions exist, we directly extract them. We then instruct $\mathbf{M}^{Q}$ to create additional questions inspired by the content of $d$. Finally, $\mathbf{M}^{Q}$ annotates each question with a special tag, either $<$is_original$>$ or $<$newly_created$>$, to indicate whether the question is an original question from $d$. The detailed prompt for $\mathbf{M}^{Q}$ is in Figure 8 of Appendix. Similar directly extraction-based synthesis method has been adopted in Yue et al. [55] Zhou et al. [56], and Yuan et al. [54].
However, the Level-1 approach described above may be limited in scalability, as the diversity and content of newly generated questions is constrained by the finite set of high-quality reference documents $\mathcal{D}$ containing questions.
Level-2 Generator.
To address the aforementioned limitation and more effectively leverage the reference document $d$, we draw inspiration from pedagogical principles. In conventional educational settings [11, 2], instructors typically begin by introducing fundamental knowledge and key concepts. They then progressively guide students toward a deeper understanding and application of these concepts through practice or exercises, thereby fostering problem-solving and critical thinking skills. Our Level-2 method mirrors this approach and implement it in two stages, as illustrated in Figure 4
@@
First, we prompt the model to act as an instructor, extracting key topics and concepts from the document $d$ for teaching. Then, we randomly sample these extracted topics and key concepts and combine them with $d$ to guide further question generation. Here, the model again acts as an instructor, creating exercises that target the sampled topics and concepts. Specifically, we prompt $\mathbf{M}^{Q}$ with a prompt $\mathbf{p}_{k}$ to extract related topics (e.g., "Curvature", "Central Limit Theorem"), and key concepts (e.g., "Arc length of a curve", "Binomial Approximation to Normal Distribution"). Formally, we compute $\mathbf{K}=\mathbf{M}^{Q}(\mathbf{p}_{k},d)$, where $\mathbf{K}$ represents the extracted meta-information described above. Note that the detailed prompt $\mathbf{p}_{k}$ is provided in Figure 11 of Appendix. Figure 12 and 13 illustrate examples of the extracted $\mathbf{K}$.
In the second stage, we focus on generating questions. To enhance the diversity of generated questions, we inject randomness into the generation process. Specifically, we first instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}$, producing a subset $\mathbf{k}_{s}$. The final question is generated by $\mathbf{M}^{Q}$, drawing on the reference document $d$, and the selected subset $\mathbf{k}_{s}$. We aim for the generated questions to remain grounded in the reference document $d$ and strictly adhere to the selected subset $\mathbf{k}_{s}$. We denote the whole question generation process above as follows:
@@
where $\mathbf{k}_{s}=\mathcal{R}(\mathbf{K})$, $\mathcal{R}$ is random selection operator and $\mathbf{x}$ is the generated questions. For simplicity, we integrate subset selection of $\mathbf{k}_{s}$ and the generation of the final questions into a single prompt $\mathbf{p}_{q}$. The detailed prompt $\mathbf{p}_{q}$ is shown in Figure 9 of Appendix. As indicated in Equation (4), the final questions $\mathbf{x}$ rely on both the (random) concept combinations generated by $\mathcal{R}(\mathbf{K})$ and the document $d$. This design intuitively offers greater diversity than the Level-1 method, which depends solely on $d$. This brings significant advantages. High-quality question synthesis is no longer constrained by the limited existing questions within reference documents. Instead, by decomposing and recombining key concepts, our Level-2 approach more efficiently exploits limited reference documents, thereby enabling more scalable synthetic question generation. In Section 5.3, we validate the benefits of Level-2 through a series of detailed ablation studies.
The overall set of topics and key concepts $\mathbf{K}$ resembles the syllabus from GLAN [24]; however, a key distinction is that each $\mathbf{K}$ in our approach is explicitly grounded in a reference document $d$.
Level-3 Generator.
Although the Level-2 generator can generate diverse questions by combining different topics and key concepts within a single document, each document inherently covers only a limited set of material. To further promote diversity of generation questions and make the approach more scalable, the Level-3 generator extends the Level-2 generator by incorporating concepts from multiple documents and grounding across multiple documents. The process of Level-3 generator is illustrated in Figure 4.
Global Concept Graph Construction.
In the Level-2 Generator, we have extracted topics and key concepts (i.e., $\mathbf{K}$) from a single document. In this section, we extend this modeling to a global perspective by considering relationships among all high-level concepts. Specifically, we construct a graph $G$, whose nodes comprise all distinct topics $\mathbb{T}$ and key concepts (KCs) $\mathbb{C}$. The intuition behind this construction is that topics or key concepts frequently co-occurring within the same document likely indicate meaningful associations. Thus, we utilize the co-occurrence statistics between topics and KCs to define the edge weights in the graph. Consequently, we consider three types of edges: topic-topic edges, topic-KC edges, and KC-KC edges, which results in three distinct sub-graphs: the topic graph, the topic-KC graph, and the KC graph. An edge between two nodes, a topic or KC $\mathbf{u}$ and another topic or KC $\mathbf{v}$, is established whenever they appear together within the same document. The edge weight $w(\mathbf{u},\mathbf{v})$ is formally defined as:
@@
where $\text{freq}(\mathbf{u},\mathbf{v})$ denotes the raw co-occurrence count of nodes $\mathbf{u}$ and $\mathbf{v}$, and $\epsilon$ is a small positive constant introduced to ensure numerical stability.
Concept Combination Sampling.
Once the concept graph $G$ has been constructed, we proceed to sample concept combinations. These sampled combinations are then utilized to retrieve relevant grounding documents, ultimately facilitating the generation of new questions.
The sampling process begins with uniform random sampling from the nodes in the topic sub-graph $\mathbb{T}$. In implementation, we iterate through all topics in $\mathbb{T}$ over multiple epochs to achieve this. Once an initial topic is randomly selected, we perform a random walk of one to two steps within the topic sub-graph to identify additional related topics. The transition probability of moving from node $\mathbf{u}$ to node $\mathbf{v}$ during the random walk is computed using a softmax function:
@@
where $\mathcal{N}(\mathbf{u})$ denotes the set of nodes adjacent to $\mathbf{u}$ in the topic sub-graph.
After obtaining the sampled topics, we next sample their corresponding key concepts. We first select an initial key concept by performing a one-step random walk on the topic-KC sub-graph, with transition probabilities computed using Equation (6) but applied specifically within this sub-graph. Subsequently, additional related key concepts are sampled through three to four steps of random walks on the KC sub-graph. This process results in a set of sampled topics and key concept $\mathbf{K}^{g}$.
Since the sampled or extracted high-level concepts essentially serve as a summary, they may omit certain details originally present in the reference documents. Thus, we search for reference documents closely aligned with our sampled concept set $\mathbf{K}^{g}$. Specifically, we calculate the Jaccard similarity between the concept set $\mathbf{K}^{g}$ and the concept sets of each document within the filtered seed document set $\mathcal{D}$. Subsequently, we select the two documents with the highest similarity scores as the final reference documents. Analogous to the procedure in the Level-2 Generator, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}^{g}$, and then utilize the selected concepts as well as the two chosen documents as references for question generation. The detailed prompt is shown in Figure 10 of Appendix. In the Level-3 Generator, the combination of concepts go beyond a single document. In addition to grounding on multiple documents, by employing random sampling on a concept graph constructed from the entire filtered document set, the diversity of the generated questions is significantly increased, thereby broadening the scope and richness of the resulting question set.
@@ 4.3 Answer Generation
Using methods described in Section 4.2, we obtain a large set questions. Now we are ready to generation answers. We employ open-source LLMs such as Qwen2.5-Math-72B-Instruct as our answer generator $\mathbf{M}^{A}$. We have not yet incorporated an additional answer verification process, since preliminary experiments indicate that our synthetic data has already yielded satisfactory results. We do believe that dedicating effort to constructing and validating answers (e.g., majority voting and multi-agent debating) can further enhance the quality of our synthetic data at additional computational cost. We plan to leave the answer verification process in further work.
@@ 5.1 Experimental Settings
Models.
For $\mathbf{M}^{Q}$, we utilize Mistral-Large-Instruct-2407 [45], as we believe that larger models inherently possess the essential internal knowledge required for effective concept extraction and question synthesis. For the answer generator, $\mathbf{M}^{A}$, we employ Qwen2.5-Math-72B [52]. For trained models, we adopt the Llama-3.2-1B, Llama-3.2-3B, and Llama-3.1-8B base models [31].
Training Settings.
We employ Supervised Fine-Tuning (SFT) to train the model. We adopt AdamW optimizer [33] and set learning rate to $1\times 10^{-5}$ with a batch size of $512$. We utilize the linear learning rate scheduler and train the model for 3 epochs.
Our Datasets.
[…]
Statistics of Synthetic Questions and Answers.
Evaluation.
To evaluate the math reasoning ability of our model trained on synthetic data, we utilize the evaluation code from Qwen-Math [52] to conduct assessments on several mainstream benchmarks: MATH [15], GSM8K [10], OlympiadBench-Math (Olympiad) [14], CollegeMath (College) [43], Minerva Math (Minerva) [22], and Gaokao 2023 En (Gaokao) [28].
Other Baselines.
We compare our synthetic data with other high-quality synthetic datasets. MAmmoTH2 [55], NaturalReasoning [54], and JiuZhang3.0 [56] employ a direct extraction-based synthesis approach on pre-training documents, similar to our Level-1 method. OpenMathInstruct-2 [47] is a large-scale synthetic math dataset consisting of 14M samples, with GSM8K and MATH serving as the seed datasets. Additionally, we compare our dataset with NuminaMath [25], a high-quality dataset that includes both synthetically generated data and extracted question-answer pairs from exam papers and mathematics discussion forums.
Furthermore, we evaluate our Level-2 and Level-3 approaches against existing methods designed to enhance question diversity through augmentation. Specifically, we include two popular augmentation approaches: the rephrase augmentation approach [53, 35] and the persona augmentation approach [12, 21], both of which generate new questions based on the questions extracted using Level-1. For rephrase augmentation, we follow the same procedure as MetaMath [53], providing two-shot examples in the prompt to rephrase questions. For persona augmentation, we utilize the 200K personas released by the authors as prompts for question augmentation.
5.2 Main Results
In this section, we present main evaluation results of SynthLLM. We test our models on various mathematical reasoning benchmarks, which encompass a range of difficulty levels. The results are shown in Table 2. The values in parentheses represent the sample size. For compared baseline training datasets, we directly evaluate model checkpoints they have released. We adopt the entire dataset synthesized with different methods from our framework to train models and denote it as SynthLLM-3B or -8B. The full dataset has a sample size of 7.4M, and we also create a subset with 3.2M samples.
SynthLLM achieves superior performance compared to other synthetic datasets.

### v3

@@ 3 SynthLLM: Web-Scale Synthetic Data Generation
In this section, we detail the SynthLLM framework for synthetic data generation. For a given target domain (e.g., mathematical reasoning), we first filter high-quality, domain-specific reference documents from Fineweb-Edu (Penedo et al., 2024), a publicly available web data repository. Using these filtered documents, we then generate large-scale, diverse questions (or prompts) by leveraging open-source LLMs through three distinct methods, each designed to progressively enhance question diversity. Finally, we generate corresponding answers to these questions, again utilizing open-source LLMs.
@@ 3.1 Reference Documents Filtering
Figure 2: Overview of reference documents filtering.
Cold-start Domain Classifier. For a given target domain, we train a binary classifier to filter web documents from Fineweb-Edu. While negative examples can be randomly sampled, obtaining sufficient positive examples for each domain is more challenging. To address this, we use synthetic documents as positive examples. Specifically, we generate synthetic domain-specific documents based on syllabi from GLAN (Li et al., 2024b), which cover 126 distinct disciplines. Each syllabus is produced by GPT-4 "(almost) from scratch", with further details available in Li et al. (2024b). As illustrated in Figure 2, we first use these syllabi to prompt a large language model (LLM) to generate synthetic documents representative of the target domain (e.g., blog posts, exam papers, and textbooks). Once we obtain the synthetic positive examples and randomly sampled negative examples from Fineweb-Edu, we train a cold-start domain classifier. Applying this classifier yields an initial set of domain-relevant documents, $\mathcal{D}^{0}$, from Fineweb-Edu.
Fine-grained Domain Classifier. The initial reference document set, $\mathcal{D}^{0}$, is typically small. To expand it, we iteratively classify Fineweb-Edu documents based on previously identified references, $\mathcal{D}^{t-1}$. At each iteration $t$, we randomly sample an additional subset $\mathcal{D}^{t-1}_{R}$ from Fineweb-Edu and instruct GPT-4o to evaluate all documents in $\mathcal{D}^{t-1}\cup\mathcal{D}^{t-1}_{R}$ on a 1–10 scale, considering relevance to the target domain, clarity, and language quality. These ratings are then used to train a fine-grained classifier, which we apply to Fineweb-Edu to construct the next iteration’s reference set, $\mathcal{D}^{t}$. We retain documents with scores above a threshold of 6.5, as it yields satisfactory performance in initial experiments. Empirically, two iterations suffice to generate a high-quality and sufficiently large reference corpus.
@@ 3.2 Document-Grounded Question Generation
After applying the filtering process outlined in Section 3.1, we obtain the refined document set $\mathcal{D}$. To generate questions, we employ three distinct methods, utilizing either single-document or multi-document contexts to progressively enhance question diversity. Let $\mathbf{M}^{Q}$ denote the LLM used for question generation.
Level-1 Generator. In the first method, we generate questions from a single reference document $d\in\mathcal{D}$. Since $d$ may already contain questions relevant to the target domain, we first prompt the LLM $\mathbf{M}^{Q}$ to determine whether $d$ includes any existing questions. If such questions exist, we directly extract them and encourage $\mathbf{M}^{Q}$ to rephrase them freely for clarity and improved understanding. We also instruct $\mathbf{M}^{Q}$ to create additional questions inspired by the content of $d$. Finally, $\mathbf{M}^{Q}$ annotates each question with a special tag, either <<original_question>> or <<newly_created>>, to indicate whether the question is an original question from $d$. The detailed prompt for $\mathbf{M}^{Q}$ is in Figure 7 of Appendix. Similar directly extraction-based synthesis method has been adopted in Yue et al. (2024), Zhou et al. (2024), and Yuan et al. (2025). However, the Level-1 approach described above may be limited in scalability, as the diversity and content of newly generated questions is constrained by the finite set of high-quality reference documents $\mathcal{D}$ containing questions.
Level-2 Generator. To overcome limitations of previous approach and better utilize $d$, we draw inspiration from pedagogical principles. In conventional education (Council et al., 2005; Anderson & Krathwohl, 2001), instructors first introduce fundamental knowledge and key concepts before guiding students toward deeper understanding through exercises, fostering problem-solving and critical thinking skills. Our Level-2 method mirrors this approach and implement it in two stages, as illustrated in Figure 3.
@@
First, we prompt the model to act as an instructor, extracting key topics and concepts from $d$ to serve as the foundation for question generation. Specifically, we prompt $\mathbf{M}^{Q}$ with $\mathbf{p}_{k}$ to identify relevant topics (e.g., "Curvature", "Central Limit Theorem") and key concepts (e.g., "Arc length of a curve", "Binomial Approximation to Normal Distribution"), formally defined as: $\mathbf{K}=\mathbf{M}^{Q}(\mathbf{p}_{k},d)$, where $\mathbf{K}$ represents the extracted meta-information. The detailed prompt $\mathbf{p}_{k}$ is provided in Figure 10, while Figure 11 and 12 illustrate examples of extracted $\mathbf{K}$. In the second stage, we generate questions while introducing controlled randomness to enhance diversity. Specifically, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}$, forming a subset $\mathbf{k}_{s}$. The final question is then generated by $\mathbf{M}^{Q}$, conditioned on both the reference document $d$ and $\mathbf{k}_{s}$. We aim for the generated questions to remain grounded in the reference document $d$ and strictly adhere to the selected subset $\mathbf{k}_{s}$. The overall question generation process is formally defined as:
@@
where $\mathbf{k}_{s}=\mathcal{R}(\mathbf{K})$, $\mathcal{R}$ is random selection operator and $\mathbf{x}$ is the generated questions. For simplicity, we integrate subset selection of $\mathbf{k}_{s}$ and the generation of the final questions into a single prompt $\mathbf{p}_{q}$, detailed in Figure 8 of the Appendix. As shown in Eq. (4), $\mathbf{x}$ rely on both (random) concept combinations generated by $\mathcal{R}(\mathbf{K})$ and $d$. This design intuitively offers greater diversity than Level-1 method, which depends solely on $d$. This brings significant advantages. High-quality question synthesis is no longer constrained by the limited existing questions within reference documents. Instead, by decomposing and recombining key concepts, our Level-2 approach more efficiently exploits limited reference documents, thereby enabling more scalable synthetic question generation. In Section 4.3, we empirically validate these benefits through a series of ablation studies. The set of topics and key concepts $\mathbf{K}$ resembles the syllabus from GLAN (Li et al., 2024b), but unlike GLAN, each $\mathbf{K}$ in our approach is explicitly grounded in a reference document $d$.
Level-3 Generator. Although Level-2 method enhances question diversity by combining different topics and key concepts within a single document, each document inherently covers only a limited scope of material. To further improve diversity and scalability, the Level-3 generator extends the Level-2 approach by incorporating concepts from multiple documents and establishing cross-document grounding. The process is illustrated in Figure 3.
Global Concept Graph Construction. In the Level-2 method, topics and key concepts (i.e., $\mathbf{K}$) are extracted from a single document. Here, we extend this modeling to a global level by capturing relationships among all high-level concepts across documents. Specifically, we construct a graph $G$, where nodes represent all distinct topics $\mathbb{T}$ and key concepts (KCs) $\mathbb{C}$. The intuition behind this construction is that topics and key concepts frequently co-occurring within the same document likely indicate meaningful associations. Thus, we define edge weights based on co-occurrence statistics. The graph consists of three types of edges: topic-topic edges, topic-KC edges, and KC-KC edges, forming three subgraphs: the topic graph, the topic-KC graph, and the KC graph. An edge is established between two nodes, $\mathbf{u}$ and $\mathbf{v}$, whenever they co-occur in the same document. The edge weight $w(\mathbf{u},\mathbf{v})$ is defined as:
@@
where $\text{freq}(\mathbf{u},\mathbf{v})$ denotes the raw co-occurrence count of $\mathbf{u}$ and $\mathbf{v}$, and $\epsilon$ is a small positive constant for numerical stability.
Concept Combination Sampling. The sampling process begins with uniform random sampling from the nodes in the topic sub-graph $\mathbb{T}$. We iterate through all topics in $\mathbb{T}$ over multiple epochs. After selecting an initial topic at random, we perform a random walk of one to two steps within the topic sub-graph to identify additional related topics. The transition probability from node $\mathbf{u}$ to node $\mathbf{v}$ is computed as:
@@
where $\mathcal{N}(\mathbf{u})$ denotes the set of adjacent nodes to $\mathbf{u}$ in the topic sub-graph. Next, we sample key concepts corresponding to the selected topics. We begin with a one-step random walk on the topic-KC sub-graph, using transition probabilities from Equation (6). Additional related key concepts are then sampled through three to four steps of random walks within the KC sub-graph, resulting in a sampled set $\mathbf{K}^{g}$. Since $\mathbf{K}^{g}$ serves as a high-level summary, certain details from the original reference documents may be omitted. Thus, we search for reference documents most aligned with $\mathbf{K}^{g}$ as grounding. Specifically, we compute the Jaccard similarity between $\mathbf{K}^{g}$ and the concept sets of each document in $\mathcal{D}$, selecting top two documents with the highest similarity scores as references. Following the Level-2 procedure, we instruct $\mathbf{M}^{Q}$ to randomly select and combine concepts from $\mathbf{K}^{g}$, using the selected concepts along with the two reference documents for question generation. The detailed prompt is provided in Figure 9 of Appendix. Unlike Level-2, which is limited to a single document, Level-3 grounds on multiple documents and employs concept sampling from a global graph, significantly enhancing the diversity and breadth of generated questions.
@@ 3.3 Answer Generation
Using the methods described in Section 3.2, we obtain a large set of questions and proceed with answer generation. We employ open-source LLMs, such as Qwen2.5-Math-72B-Instruct, as our answer generator $\mathbf{M}^{A}$. Currently, we do not incorporate an additional answer verification process, as preliminary experiments indicate that our synthetic data already achieves satisfactory quality. The detailed experimental results are shown in Section C.3. However, further improvements could be made by refining answer validation through techniques like majority voting (Wang et al., 2022; Jiao et al., 2024) or multi-agent debate (Du et al., 2024; Khan et al., 2024), albeit at additional inference computational cost. We leave this for future work.
@@ 4.1 Experimental Settings
Models and Our Datasets. For $\mathbf{M}^{Q}$, we use Mistral-Large-Instruct-2407 (team, 2024), as larger models inherently possess the internal knowledge necessary for effective concept extraction and question generation. For $\mathbf{M}^{A}$, we employ Qwen2.5-Math-72B (Yang et al., 2024). The trained models are Llama-3.2-1B, -3B, and Llama-3.1-8B base models (Llama Team, 2024). After filtering, we collected approximately $520\text{K}$ math documents. For Level-1 and Level-2, we generated 5 questions per reference document, yielding approximately $2.3\text{M}$ and $2.6\text{M}$ samples, respectively. For Level-3, each sampled $\mathbf{K}^{g}$ generated 3 questions, resulting in $2.6\text{M}$ samples. Further details of training and dataset are provided in Section B.
Other Baselines and Evaluation Benchmarks. We compare our approach with MAmmoTH2 (Yue et al., 2024), JiuZhang3.0 (Zhou et al., 2024), NaturalReasoning (Yuan et al., 2025), and OpenMathInstruct-2 (Toshniwal et al., 2025). Additionally, we evaluate our Level-2 and Level-3 methods against existing augmentation techniques aimed at enhancing question diversity. Specifically, we consider two widely used approaches: rephrase augmentation (Yu et al., 2024) and persona augmentation (Ge et al., 2024; Lambert et al., 2024; Wang et al., 2025), both of which generate new questions based on those extracted in Level-1. Following Qwen-Math (Yang et al., 2024), we adopt evaluation benchmark: MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), OlympiadBench-Math (Olympiad) (He et al., 2024), CollegeMath (College) (Tang et al., ), Minerva Math (Minerva) (Lewkowycz et al., 2022), and Gaokao 2023 En (Gaokao) (Liao et al., 2024). The more details are shown in Section B.
4.2 Main Experimental Results
We evaluate SynthLLM across multiple mathematical reasoning benchmarks of varying difficulty levels. Table 2 presents results, with values in parentheses indicating sample sizes. For baseline datasets, we directly evaluate their released models. We train models using both the full SynthLLM dataset (7.4M samples) and a subset obtained by uniformly sampling 3.2M examples, denoted as SynthLLM-1B, SynthLLM-3B, and SynthLLM-8B.